# HOMURA: Taming the Sand-Glass for Time-Constrained LLM Translation via Reinforcement Learning

### v1

@@ Appendix H Examples of Structured CoT Reasoning for GenRM
Appendix M Theoretical and Empirical Analysis of Regularization Strength
In Eq.(4), we defined the KL-regularized objective where $\beta$ controls the trade-off between control strength and faithfulness to the base model $P_{0}$. In our experiments, we observed that setting $\beta\to 0$ leads to superior syllable-ratio compliance and linguistic quality. This appendix provides a rigorous justification for why the stability typically guaranteed by explicit KL regularization is preserved in HOMURA through the GRPO framework, even in the control-only regime ($\beta=0$).
M.1 Theoretical Justification: From Token-level to Sequence-level Regularization
Standard RLHF objectives incorporate a token-level KL penalty to anchor the policy to a reference model. As shown in Eq.(4), this term imposes a local constraint on every transition, forcing $\pi_{\theta}$ to minimize divergence from the prior of $P_{0}$ at each step. For time-constrained translation (Sand-Glass), satisfying the control condition $c$ requires radical lexical pruning and syntactic restructuring. Under these conditions, the KL term acts as high-dimensional friction, making structural deviations mathematically expensive and hindering the model from reaching the optimal rate-distortion frontier.
By setting $\beta=0$, we prioritize control strength over faithfulness, granting the model maximum compression freedom. Numerical stability is maintained through the synergy of two stabilizing forces:
  • Algorithmic Stability via Centered Clipped Objective: Stability is maintained through a surrogate loss function $\mathcal{L}_{GRPO}$ that integrates group-relative normalization with a clipping mechanism. The policy update is derived from:
@@
    where $r_{i,t}(\theta)$ is the token-level probability ratio, and the advantage $\hat{A}_{i}$ is normalized within the group:
    $\hat{A}_{i}=\frac{R(x,c,y_{i})-\mu_{group}}{\varsigma_{group}},\quad\mu_{group}=\frac{1}{G}\sum_{j=1}^{G}R(x,c,y_{j})$ (17)
    This formulation ensures stability via (1) Relative Anchoring, where the zero-sum nature ($\sum\hat{A}_{i}=0$) prevents global policy drift by centering updates around the group’s moving average; and (2) Trust Region Enforcement, where clipping mitigates policy staleness [39]. Together, they preserve the validity of the first-order approximation for sequence-level rewards without requiring an explicit KL penalty.
  • Implicit Structural Regularization: The syllable-ratio reward $R_{\rho}$ itself functions as a structural regulator. By penalizing linguistic redundancy, $R_{\rho}$ induces a form of structural sparsity, forcing the model to operate within a lower-dimensional manifold of the total policy space. This task-specific constraint effectively prunes suboptimal trajectories that are semantically valid but temporally excessive, providing an alternative stabilization force that aligns the optimization path directly with the Sand-Glass objective.
M.2 Empirical Analysis
To validate these theoretical insights, we analyze the training process across different KL coefficients (Figure 7).
Figure 7: Training dynamics across $\beta\in\{0,0.01,0.05\}$. From left to right: (1) KL Loss, (2) Entropy Loss, (3) Policy Gradient Loss, (4) Gradient Norm, and (5) Syllable Ratio. The $\beta=0$ variant (light blue) achieves the best balance between exploration and stability.
Dynamics and Gradient Vitality.
As illustrated in Figure 7, the $\beta=0$ regime (light blue) demonstrates superior optimization vitality. By removing the faithfulness constraint, the model maintains a healthy and stable Gradient Norm ($\sim$4.0) and robust Policy Gradient Loss, confirming that reward signals are effectively translated into policy updates without being neutralized by optimization friction.
Notably, while $\beta=0$ maintains a high and purposeful Policy Entropy (signifying broad compression-driven exploration), the $\beta=0.05$ variant (purple) exhibits an abnormal surge in entropy mid-training. We interpret this as optimization paralysis: the excessive KL penalty forces the probability distribution to flatten as the model becomes indecisive between length compliance and faithfulness to the verbose prior. This conflict leads to a stagnating Gradient Norm in $\beta>0$ variants, where the model settles for conservative, localized edits once it hits the KL-constrained boundary.
Quantitative Performance: Non-linear Trade-offs.
Table 12 highlights a non-linear relationship between KL regularization and performance. While all configurations successfully satisfy the syllable budget, their linguistic strategies differ:
Table 12: Impact of $\beta$ on final performance. Reaching the budget through $\beta=0$ preserves significantly more linguistic structure (CERR) by favoring structural adaptation over conservative edits. IB denotes the inclusion of the Information Bottleneck module.
@@
  • The Superiority of $\beta=0$: The $\beta=0$ variant achieves the highest CERR and BLEU-$\rho$, significantly outperforming all regularized settings. This confirms that total liberation from the token-level prior is essential for structural adaptation—the model’s ability to fundamentally rewrite sentences for conciseness rather than merely deleting tokens.
  • The Conservative Trap of $\beta>0$: Interestingly, while $\beta=0.05$ shows a slight recovery in COMET compared to $\beta=0.01$, its CERR remains substantially lower than the $\beta=0$ baseline. This suggests that higher $\beta$ values force the model into a conservative pruning regime: it clings to the lexical surface of the reference to satisfy the KL penalty (thereby maintaining COMET levels), but at the cost of failing to find the optimal structural alignment (low CERR and BLEU-$\rho$), eventually damaging the natural flow of the synchronized output.
The data confirms that the lack of optimization momentum (decaying Gradient Norm in Figure 7) for $\beta>0$ prevents the model from escaping the reference model’s verbosity trap. By setting $\beta=0$, the policy can effectively navigate the complex rate-distortion frontier required for high-precision synchronization.

### v2

@@ Appendix H Examples of Structured CoT Reasoning for GenRM
Appendix M Analysis of Regularization Strength
In Eq.(4), $\beta$ controls the trade-off between control strength and adherence to the reference model $P_{0}$. Our experiments show that $\beta\to 0$ leads to better syllable-ratio compliance and linguistic quality. This appendix justifies how HOMURA maintains training stability and semantic integrity even without explicit KL regularization ($\beta=0$).
M.1 Theoretical Justification: Stability through Implicit Regularization
A common concern in RLHF is that removing the token-level KL penalty ($\beta=0$) may lead to training instability. However, we argue that for time-constrained translation, the explicit KL penalty defined in Eq. (4) is not only detrimental but also redundant, as its stabilizing role is effectively superseded by the internal dynamics of our framework.
Escaping the Verbosity Trap.
Standard RLHF uses token-level KL divergence to anchor the policy to a pre-trained reference $P_{0}$. In the Sand-Glass task, the model must perform significant structural compression to meet strict temporal constraints. A rigid KL penalty treats these necessary structural deviations as distributional shifts, mathematically penalizing the model for being concise and confining it to the verbose distribution of the base model. By setting $\beta=0$, we transition from distribution matching to goal-oriented optimization, allowing the model to prioritize control strength and efficiency.
Sufficiency of Implicit Regularization.
The stability typically provided by the KL term is maintained through two alternative forces:
  • Token-level Trust Region (GRPO Dynamics): Numerical stability is preserved via the surrogate loss function $\mathcal{L}_{GRPO}$, which integrates two complementary token-level stabilizing forces:
@@
    where the advantage $\hat{A}_{i}$ is normalized within the group: $\hat{A}_{i}=(R_{i}-\mu_{\text{group}})/\varsigma_{\text{group}}$.
    The clipping mechanism enforces a trust region relative to the evolving policy $\pi_{\theta_{old}}$, preventing abrupt updates. Simultaneously, the group-relative normalization centers updates around the group’s moving average. This relative anchoring ensures the policy explores the length-constrained space without global drift.
  • Sequence-level Implicit Semantic Regularization ($R_{bt}$): While the token-level KL constraint is removed, the policy is still constrained at the sequence level by the structural-agnostic back-translation reward $R_{bt}$. This acts as an implicit semantic regularizer: it does not penalize how the model expresses information (lexical choice), but rather whether the information is preserved (semantic fidelity). This sequence-level anchoring allows the model to explore radical syntactic restructuring for conciseness while ensuring the output remains a faithful translation.
In summary, the synergy between GRPO’s token-level trust region and the sequence-level semantic anchor of $R_{bt}$ provides a sufficient regularizing effect. This allows HOMURA to escape the reference’s verbosity trap and navigate the complex rate-distortion frontier required for high-precision synchronization.
M.1 Empirical Analysis
We analyze the training dynamics across different KL coefficients $\beta\in\{0,0.01,0.05\}$ in Figure 7.
Figure 7: Training dynamics for varying $\beta$. The $\beta=0$ variant (light blue) maintains a stable gradient norm and policy entropy, indicating effective exploration without optimization stagnation.
Training Dynamics and Gradient Norms.
As shown in Figure 7, the $\beta=0$ regime (light blue) demonstrates superior optimization vitality. Without the KL constraint, the model maintains a stable Gradient Norm ($\sim$4.0) and a healthy Policy Gradient Loss. In contrast, the $\beta=0.05$ variant (purple) shows a significant increase in policy entropy mid-training followed by a stagnating gradient norm. This suggests that a high KL penalty creates conflicting objectives between length compliance and reference faithfulness, leading to inefficient policy updates where the model fails to find a clear optimization direction.
Performance Trade-offs: The Pruning vs. Rewriting.
Table 12 shows the impact of $\beta$ on final metrics. While all configurations achieve the target syllable budget, their linguistic strategies differ significantly:
Table 12: Impact of $\beta$ on performance. $\beta=0$ achieves the best balance between compression and fidelity, as indicated by BT-CERR and BLEU-$\rho$.
@@
The $\beta=0$ configuration yields the highest BT-CERR and BLEU-$\rho$, outperforming all regularized settings. This indicates that removing the token-level prior is essential for structural adaptation—the model’s ability to fundamentally rewrite sentences for conciseness rather than just deleting words. Higher $\beta$ values force the model into a conservative strategy: it adheres closely to the reference’s lexical surface to satisfy the KL penalty (maintaining higher COMET), but fails to achieve optimal structural efficiency (lower BT-CERR), which is critical for high-precision synchronization in time-constrained scenarios.